OpenStudy (anonymous):

im really struggling with translating word problems into equations. "Andy as 19 coins made up of quarters and half dollars, their total value i $5.25. How many quarters does he have?" so the numbers i have to work with are 19, 25, 50 and 5.25. Im not sure how to arrange them properly

OpenStudy (irishboy123):

maybe start with the idea that: Q = no of quarters H = no of half dollars ie; the actual number of coins. so \(19 = Q + H \qquad🐞\) make sense? see the second equation ?

OpenStudy (anonymous):

Wow YEP xD

OpenStudy (irishboy123):

further hint: so if he has Q quarters in his purse, he has $(\(Q \times \frac{1}{4}\)) in there.

OpenStudy (anonymous):

I'm... still kind of confused.. I think... I'm not sure if I have this right-- would it be: Q + H = 19 25Q + 50H = 5.25? and then solve by elmination ?

jimthompson5910 (jim_thompson5910):

@SmileyVizard The first equation you have is correct, but the second `25Q + 50H = 5.25` is not correct

jimthompson5910 (jim_thompson5910):

25Q represents the total value of just the quarters in CENTS (not dollars) same with 50H so you have to convert $5.25 to cents 5.25 dollars = 5.25*100 = 525 cents

jimthompson5910 (jim_thompson5910):

So the second equation would be `25Q + 50H = 525` or you can convert everything to dollars to get `0.25Q + 0.50H = 5.25`

OpenStudy (anonymous):

so the equation as a whole would be Q + H = 19 25Q + 50H = 525?

jimthompson5910 (jim_thompson5910):

yes that is the system of equations

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

do you see how to find Q and H from here?

OpenStudy (anonymous):

yeah i think so

jimthompson5910 (jim_thompson5910):

ok great, tell me what you get

OpenStudy (anonymous):

okay-- give me a minute to solve it

jimthompson5910 (jim_thompson5910):

sure thing, take all the time you need

OpenStudy (anonymous):

i got H = 2 and Q = 17

jimthompson5910 (jim_thompson5910):

you have the correct answer check: Q+H = 19 17+2 = 19 19 = 19 ... works 25Q + 50H = 525 25*17 + 50*2 = 525 525 = 525 ... works solution (Q,H) = (17,2) is confirmed so this proves she has 17 quarters and 2 half dollars

OpenStudy (anonymous):

yes!!! thank you so much for your help!

jimthompson5910 (jim_thompson5910):

you're welcome